whats four times the sum of three consecutive integers is 156. what are the integers?
Answer by checkley71(8403)

LET X BE THE FIRST INTEGER.
LET X+1 BE THE NEXT INTEGER.
LET X+2 BE THE THIRD INTEGER.
NOW 4 TIMES THE SUM OF THESE 3 INTEGERS=156.
HOPE THIS HELPS. : )
-----------------------------------------
4(X+[X+1]+[X+2])=156
4(3X+3)=156
12X+12=156
12X=156-12
12X=144
X=144/12
X=12 FOR THE FIRST INTEGER.
12+1=13
12+2=14
PROOF:
4(12+13+14)=156
4*39=156
156=156